OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (anonymous):

a=-9 d=-3-(-9)=-3+9=6 l=a+(n-1)d 81=-9+(n-1)6=-9+6n-6=6n-15 6n=81+15=96 n=96/6=16 so no.of terms=16 \[S _{n}=\sum_{n=1}^{16}\left( 6n-15 \right)\]

OpenStudy (anonymous):

it is also correct.